### v1

@@ A.2 Needle-in-a-Haystack
Figure 11: Needle-in-a-Haystack results of LLaMA-3.2-3B-Instruct.
@@
A.3 Latency and Throughput Evaluation
We present the TTFT and the throughput results of the baselines and FastKV in Table 5 and Table  7. The throughput results of Full KV is depicted in Table  6 for effective comparison. The throughput results of KV cache compression methods are evaluated with input context length 128k. As the input context length increases, the relative TTFT gain of FastKV compared to SnapKV, AdaKV and HeadKV increases. For Mistral-Nemo-12B-instruct, FastKV achieves TTFT similar to GemFilter because the TSP layer index is same with the GemFilter filter layer. The token throughput gain of the baselines and FastKV using fixed KV budget grows rapidly as the throughput of Full KV declines due to the increasing size of KV cache. FastKV consistently achieves the highest throughput across all scenarios.
@@ A.3 Latency and Throughput Evaluation
Method 8k 16k 32k 64k 128k
LLaMA-3.1-8B-Instruct
     Full KV 0.64 1.41 3.40 9.05 27.94
     SnapKV 0.67 1.45 3.49 9.22 28.27
     AdaKV 0.75 1.53 3.57 9.31 28.36
     HeadKV 0.70 1.48 3.51 9.25 28.30
     GemFilter 0.44 0.78 1.66 4.15 12.51
     FastKV 0.41 0.80 1.81 4.66 14.15
@@
     Full KV 0.92 2.04 4.85 12.71 37.94
     SnapKV 0.95 2.08 4.90 12.80 38.11
     AdaKV 1.05 2.18 5.01 12.92 38.23
     HeadKV 0.99 2.12 4.95 12.85 38.18
     GemFilter 0.69 1.25 2.65 6.59 19.22
     FastKV 0.59 1.16 2.57 6.53 19.22

### v2

@@ C.2 Needle-in-a-Haystack
C.3 Latency and Throughput Evaluation
We present the TTFT and throughput results of baseline methods and FastKV in Table 6 and Table 8. The TTFT of LLaMA-3.1-8B-UltraLong-1M-Instruct is measured with input context lengths up to 256K tokens, while for Mistral-Nemo-12B-Instruct, the measurement is limited to 128K tokens due to its 128K context window size. The throughput results for Full KV are provided separately in Table 7 for effective comparison. The throughput results for KV cache compression methods are evaluated at an input context length of 128K tokens.
As the input context length increases, the relative TTFT gain of FastKV over SnapKV, AdaKV, and HeadKV becomes more pronounced. For Mistral-Nemo-12B-Instruct, FastKV achieves TTFT comparable to GemFilter, as both utilize the same TSP layer structure as the GemFilter filtering layer. The token throughput of the baselines and FastKV with a fixed KV budget improves rapidly as the throughput of Full KV declines with increasing KV cache size. FastKV consistently achieves the highest throughput across all scenarios.
@@ C.3 Latency and Throughput Evaluation
      Method       8K       16K       32K       64K       128K       256K
      LLaMA-3.1-8B-UltraLong-1M-Instruct
      Full KV       0.64       1.41       3.40       9.05       29.2       OOM
      SnapKV       0.67       1.45       3.49       9.22       29.68       102.60
      AdaKV       0.75       1.53       3.57       9.31       29.88       103.12
      HeadKV       0.70       1.48       3.51       9.25       29.65       102.71
      GemFilter       0.44       0.80       1.66       4.15       12.95       45.08
      FastKV       0.41       0.80       1.81       4.66       14.80       51.68
@@
      Full KV       0.92       2.04       4.85       12.71       37.94       -
      SnapKV       0.95       2.08       4.90       12.80       38.11       -
      AdaKV       1.05       2.18       5.01       12.92       38.23       -
      HeadKV       1.0       2.1       5.0       12.9       38.18       -
      GemFilter       0.69       1.25       2.65       6.59       19.22       -
      FastKV       0.59       1.16       2.57       6.53       19.22       -